OpenStudy (anonymous):

ln56/ln7=ln8. Is this true or is this false?

OpenStudy (anonymous):

When I inputted them into my calculator. ln8 was not exactly as the same answer ln56/ln7 but it was very close. Would this make it true nonetheless? ln56/ln7=4.025/1.946=2.069. but ln8=2.047

OpenStudy (anonymous):

According to my calculator, it is false.

OpenStudy (anonymous):

Did you get different values? if I use log I get a different answer as well

OpenStudy (anonymous):

When you divide ln's, maintain the function and divide the terms. Hence ln56/ln7=ln(56/7)=ln8, therefore, true.

OpenStudy (anonymous):

Interesting. Then 56/7 can be divided when next to ln?

OpenStudy (anonymous):

My calculator has In sign on so that's i got it's false.

OpenStudy (anonymous):

Wait, wait!! It is false!

OpenStudy (anonymous):

Please pay attention the following properties:

OpenStudy (anonymous):

\[lnA-lnB=\ln (\frac{ A }{ B })\] \[lnA:lnB=\ln(A-B)\] \[lnA*lnB=\ln(A+B)\]\[lnA+lnB=\ln(A*B)\] Hence\[\ln56/\ln7=\ln(56-7)=\ln49\neq \ln8\]

OpenStudy (anonymous):

Ah, I see. I forgot to implement the logarithm rules. Thank you for very much. This is perfect

OpenStudy (anonymous):

No problem.